A NaOH solution is standardized using axalic acid, C2H2O4 (90g/mol) according to the following balanced equation: H2C2O4 (aq) + 2NaOH (aq) -> Na2C2O4 (aq) + 2H2O (l)

If 0.8986g of oxalic acid requires 63.92mL of NaOH, what is the NaOH concentration?
Answer #1

Balanced chemical equation is:

H2C2O4 (aq) + 2NaOH (aq) -> Na2C2O4 (aq) + 2H2O (l)

Molar mass of H2C2O4,

MM = 2*MM(H) + 2*MM(C) + 4*MM(O)

= 2*1.008 + 2*12.01 + 4*16.0

= 90.036 g/mol

mass(H2C2O4)= 0.8986 g

use:

number of mol of H2C2O4,

n = mass of H2C2O4/molar mass of H2C2O4

=(0.8986 g)/(90.04 g/mol)

= 9.98*10^-3 mol

According to balanced equation

mol of NaOH reacted = (2/1)* moles of H2C2O4

= (2/1)*9.98*10^-3

= 1.996*10^-2 mol

This is number of moles of NaOH

volume , V = 63.92 mL

= 6.392*10^-2 L

use:

Molarity,

M = number of mol / volume in L

= 1.996*10^-2/6.392*10^-2

= 0.3123 M

Answer: 0.312 M